OpenStudy (anonymous):

Factors of 5000x^3 + 5?

OpenStudy (mertsj):

\[5000x^3+5=5(1000x^3+1)=5(10x+1)(100x^2-10x+1)\]

OpenStudy (campbell_st):

5(1000x^3 +1) this is now the sum of 2 cubics inside the brackets0 let a = 10x and b = 1 then a^3 + b^3 = (a +b)(a^2 -ab + b^2)

OpenStudy (anonymous):

Thank you :)

OpenStudy (mertsj):

yw